OpenStudy (anonymous):

The product of two consecutive odd integers is 1 less than 6 times their sum. Find the two integers.

OpenStudy (anonymous):

let a is any odd integer, a*(a+2)=6*(a+a+2)-1 \[a ^{2}-10a-11=0\] \[a ^{2}+a-11a-11=0\] \[a(a+1)-11(a+1)=0\] \[(a+1)(a-11)=0\] a=-1,11

OpenStudy (anonymous):

my online work said -1 was correct, but I need 2 different sets of values for my problem.

OpenStudy (anonymous):

@KingD @tanya123 , Are y'all going to help oooooor...?

OpenStudy (tanya123):

Sorry but I dont know!

OpenStudy (tanya123):

and I think @hanzi has already helped you.../

OpenStudy (anonymous):

I need more answer, plus one was incorrect :/

OpenStudy (anonymous):

-1 and 11 are the answers

OpenStudy (anonymous):

@ganeshie8 , please help :(

OpenStudy (anonymous):

I need 2 more answers, I'll take a picture...

OpenStudy (anonymous):

-1 and (-1+2)=1 so, one set of answers is -1,1 similarly other set is 11,(11+2) 11,13

OpenStudy (tkhunny):

Note: a and a+2 are not necessarily odd. If you get a = 2, you have a problem. If you REALLY want odd, perhaps 2k+1, 2k+3 for k an integer. This may be better, as it truly defines odd values, but we are still relying a little on luck. What if k isn't an integer in our final solution.

OpenStudy (anonymous):

Got it! Thank you ! :)